OpenStudy (anonymous):

Write an equation in slope-intercept form of the line through point P(6, 7) with slope –2.

OpenStudy (anonymous):

A y – 7 = –2(x – 6) B y = –2x + 7 C y = –2x + 19 D y – 6 = –2(x – 7)

OpenStudy (anonymous):

in point slope form, y-y1=m(x-x1) plug in those values, then solve for y

OpenStudy (anonymous):

m=-2 x1=7 y1=7

OpenStudy (anonymous):

soo it would be y-7=-2(x-6) ?

OpenStudy (anonymous):

that's point-slope form, not what the question wanted. just solve for y to get slope-intercept.

OpenStudy (anonymous):

how do i solve for y?